OpenStudy (anonymous):

solve to the nearest degree for all the values of Ɵ in the interval 0≤Ɵ≤360 2cos^2Ɵ-4cosƟ+1=0

OpenStudy (anonymous):

First we need to use substitution for this like any variable for cos Lets assume that \[\cos \theta=m\] So we get equation as \[2m^2-4m+1=0\] Getting this ???

OpenStudy (anonymous):

i get that but i think im getting the wrong answer

OpenStudy (anonymous):

Have you solve this quadratic equation ??

OpenStudy (anonymous):

yes but im not sure how to proceed after i get what m equals to

OpenStudy (anonymous):

nevermind i got it

OpenStudy (anonymous):

okay ... byeee :)